How would you rationalize the denominator of this expression: square root of x divided by square root of 8y
Answer by MathPro(14)

Good Afternoon,
Your problem is:
How would you rationalize the denominator of this expression: square root of x divided by square root of 8y
I’ll write that down using the following notation:
Square root x = sqrt(x) so your problem looks like:
sqrt(x) / sqrt(8y)
This problem can be solved a lot easier if we factor the denominator first:
The number 8 can be rewritten as 2 (2)(2) so we have 2 squared times 2 Lets rewrite the problem now as
sqrt(x)/ sqrt[(2(2)(2)y]=
sqrt(x)/[sqrt[2(2)] sqrt(2y)]
note that sqrt[2(2)] = 2 so we can rewrite the problem as
sqrt(x) / [2sqrt(2y)]
We want to get rid of the square root on the bottom to rationalize it. The product rule states that:
sqrt(x) sqrt(x) = x ^ (1/2 + 1/2) = x^1 = x, where ^ means raise to the power
To apply the product rule to our problem multiply the bottom and top by sqrt(2y) to get:
[ sqrt(2y)sqrt( x)] / [2 sqrt(2y)sqrt(2y)] =
sqrt(2yx) / [2 (2y)] =
sqrt(2yx) / (4y)
Good Luck!
Selma from Rent-A-Whiz, an online tutoring and homework help website. Just visit www.rentawhiz.com to learn more. Currently, we are offering to solve one math problem for free. Just visit our website, select the homework tab, then click on the math problem submission link and quote coupon code 003 when you submit your problem.